RECAPITALIZATION Tartan Industries currently has total capital equal to $5 million, has zero debt, is in the 40% federal-plus-state tax bracket, has a net income of $1 million, and distributes 40% of its earnings as dividends. Net income is expected to grow at a constant rate of 4% per year, 280,000 shares of stock are outstanding, and the current WACC is 12.10%. The company is considering a recapitalization where it will issue $3 million in debt and use the proceeds to repurchase stock. Investment bankers have estimated that if the company goes through with the recapitalization, its before-tax cost of debt will be 11% and its cost of equity will rise to 16.5%. a. What is the stocks current price per share (before the recapitalization)? Round your answer to the nearest cent. Do not round intermediate steps b. Assuming that the company maintains the same payout ratio, what will be its stock price following the recapitalization Assume that shares are repurchased at the price calculated in part a. Round your answer to the nearest cent. Do not round intermediate steps.
Answer #1

a) All equity invested with capital of $5 million

Current price per share (before recapitalisation) = Dividend per share (DPS) / (Cost of equity - Growth rate)

Here, Cost of equity (ie WACC) # = 12.10% or 0.1210 ; Growth rate = 4% or 0.04

DPS = Total dividend to equity shareholders / No. Of equity shares

Dividend to equity shareholders = Net income *Dividend payout ratio = $1 million * 0.40 = $0.40 million or 4,00,000

No. Of equity share = 2,80,000 shares

DPS = $4,00,000 / 2,80,000 shares = $1.4286 per share

Now,

Current price per share (before recapitalisation) = $1.4286 / ( 0.1210 - 0.04) = $1.4286 / 0.081 = $17.64 per share

b) Price of equity per share (after recapitalisation) = DPS /(Cost of equity - Growth rate)

Here, Cost of equity = 16.5% or 0.1650 ; Growth rate = 4% or 0.04

DPS (after recapitalisation) = Dividend to equity shareholders / No. Of equity shares

i) Dividend to equity shareholders = (Net ncome - (Interest on debt * (1 -tax rate))) * Dividend payout ratio

here, Net income = $1 million

Interest on debt = $3 million * 11% = $0.33 million

Dividend payout ratio = 40%

We will put the values into the formula

Dividend to equity shareholders = ($1 million - (0.33 million * (1 - 0.40)) * 0.40 = ($1 - ($0.33 * 0.60)) * 0.40

Dividend to equity shareholders = ($1 - $0.1980) * 0.40 = $0.8020 * 0.40 = $0.3208 million or $3,20,800

ii) No. Of equity shares (after recapitalisation) = Existing equity shares - (repurchased equity shares)

Here,

Repurchased equity shares = Repurchased amount / Equity price before recapitalisation

Repurchased equity shares = $3 million or $30,00,000 / $17.64 = 1,70,068 shares

No. Of equity shares (after recapitalisation) = 2,80,000 shares - 1,70,068 shares = 1,09,932 shares

We will put the values into DPS formula

DPS = $3,20,800 / 1,09,932 shares = $2.9182 per share

Now,

Price of equity per share (after recapitalisation) = $2.9182 per share / (0.1650 - 0.04) = $2.9182 per share / 0.125

Price of equity per share (after recapitalisation) = $23.35 per share